OpenStudy (anonymous):

At a certain factory, products are made in vats that have the capacity to hold 20 quarts. If each bottle of the product contains 4/115 quarts, how many bottles can be made from each vat?

OpenStudy (anonymous):

i'm not sure: since in 1 vat=20quarts and 1 bottle=4/11 quarts 20/(4/11)= 55 bottles.. am i right? @agentsmith

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Divide 20 by 4/115

OpenStudy (anonymous):

nice new pic @agent0smith lol ummm so i divide 4/115 right 4 into 115

OpenStudy (anonymous):

so is it even 55 ?

OpenStudy (anonymous):

so what i did is right? (^_^)

OpenStudy (anonymous):

but u put 11 not 115 lol

OpenStudy (anonymous):

OH sweet!!! my bad :/ so it's not 55..

OpenStudy (anonymous):

so how would i divide 20 by 4/115 do i divide the fraction first

OpenStudy (agent0smith):

When dividing by a fraction, flip it and multiply by the reciprocal... \[\Large \frac{ 20 }{ 1 } \div \frac{ 4 }{ 115} = \frac{ 20 }{ 1 } \times \frac{ 115 }{ 4 }\]

OpenStudy (anonymous):

575 bottles

OpenStudy (anonymous):

You are correct my dear sir! Have yourself a wonderful day!